### v1

# SPAN: Benchmarking and Improving Cross-Calendar Temporal Reasoning of Large Language Models

Abstract
Temporal reasoning is a fundamental capability for large language models (LLMs) to understand real-world dynamics. Existing research on temporal reasoning has predominantly focused on the Gregorian calendar. However, as many countries and regions concurrently adopt multiple calendar systems, temporal reasoning across calendars becomes crucial for LLMs in global and multicultural contexts. Unfortunately, cross-calendar temporal reasoning remains underexplored, with no dedicated benchmark available to evaluate this capability. To bridge this gap, we introduce SPAN, a croSs-calendar temPoral reAsoning beNchmark, which requires LLMs to perform intra-calendar temporal reasoning and inter-calendar temporal conversion. SPAN features ten cross-calendar temporal reasoning directions, two reasoning types, and two question formats across six calendars. To enable time-variant and contamination-free evaluation, we propose a template-driven protocol for dynamic instance generation that enables assessment on a user-specified Gregorian date. We conduct extensive experiments on both open- and closed-source state-of-the-art (SOTA) LLMs over a range of dates spanning $100$ years from $1960$ to $2060$. Our evaluations show that these LLMs achieve an average accuracy of only 34.534.5%, with none exceeding 8080%, indicating that this task remains challenging. Through in-depth analysis of reasoning types, question formats, and temporal reasoning directions, we identify two key obstacles for LLMs: Future-Date Degradation and Calendar Asymmetry Bias. To strengthen LLMs’ cross-calendar temporal reasoning capability, we further develop an LLM-powered Time Agent that leverages tool-augmented code generation. Empirical results show that Time Agent achieves an average accuracy of 95.3195.31%, outperforming several competitive baselines, highlighting the potential of tool-augmented code generation to advance cross-calendar temporal reasoning. We hope this work will inspire further efforts toward more temporally and culturally adaptive LLMs. Our source code and datasets are available at https://github.com/miaozhongjian/span.git.
Introduction
Recent advances in large language models (LLMs) have led to substantial progress across a range of reasoning tasks (qiao-etal-2023-reasoning; xu2024faithful; li2024understanding; fan-etal-2025-slam). Among these, temporal reasoning, as a fundamental capability for LLMs to understand real-world dynamics, has garnered increasing attention, leading to the development of multiple evaluation benchmarks (chen2021datasetansweringtimesensitivequestions; liska2022streamingqa; kasai2023realtime; tan2023benchmarkingimprovingtemporalreasoning; wei-etal-2023-menatqa; tan-etal-2023-towards; fatemi2024test; wang2024trambenchmarkingtemporalreasoning; chu-etal-2024-timebench; wang-zhao-2024-tram; ge2025tremu; wei2025time; saxena2025lost). Nevertheless, most existing efforts focus on temporal reasoning within the Gregorian calendar, neglecting the exploration of LLMs’ capability to reason across calendar systems. In practice, numerous countries and regions employ an alternative calendar system alongside the widely-used Gregorian calendar system (taqizadeh1939various; graumann2015problem). For instance, both Saudi Arabia and China use the Gregorian calendar for official purposes, but Saudi Arabia follows the Islamic calendar for religious observances while China relies on the Chinese lunar calendar for traditional cultural activities. These real-world practices underscore the necessity for reasoning and converting dates across calendar systems, i.e., cross-calendar temporal reasoning, which is crucial for global and multicultural applications of LLMs.
Despite its practical significance, the cross-calendar temporal reasoning capability of LLMs remains unexplored. Critically, there is currently no benchmark for assessing this capability, which hinders further research progress. Moreover, existing temporal reasoning benchmarks are inherently time-invariant, limiting evaluation to a single, fixed reference date when the benchmark was created. This leads to two drawbacks: (1) Temporal Scope Limitation. Given the dynamic nature of time, robust temporal reasoning evaluation should ensure a broad temporal scope by enabling assessments at arbitrary reference dates, which is not yet supported by existing benchmarks. (2) Data Contamination. The ground truth in certain temporal reasoning benchmarks is typically deterministic and publicly available through platforms such as Hugging Face and GitHub. Such accessibility may lead to evaluation data being absorbed into training corpora, potentially biasing evaluations (sainz-etal-2023-nlp; deng-etal-2024-investigating; balloccu2024leak).
In response to these challenges, we introduce SPAN, a croSs-calendar temPoral reAsoning beNchmark. Unlike prior Gregorian-centric temporal reasoning benchmarks, SPAN requires LLMs to perform temporal reasoning within one calendar based on a reference date, and then convert the result to its equivalent in another calendar. To mitigate data contamination and time-invariant problems, we propose a novel evaluation protocol that dynamically generates instances by instantiating a set of question–code template pairs. By utilizing these carefully-designed template pairs, SPAN generates diverse evaluation instances covering three key dimensions: (1) Temporal Reasoning Directions. SPAN supports ten cross-calendar reasoning directions between the Gregorian and five other calendars: Chinese Lunar, Shaka, Hebrew, Islamic, and Persian. (2) Reasoning Types. SPAN encompasses date- and festival-based reasoning tasks that perform cross-calendar conversions for general dates and festival dates, respectively. (3) Question Formats. SPAN comprises both polar and content questions, where the answers to polar questions are binary judgments (i.e., “Yes” or “No”), whereas content questions require calendar-specific dates as answers. During evaluation, these question-code template pairs are instantiated to generate questions and corresponding code snippets, which are then executed to automatically derive the answers. Overall, SPAN offers a dynamic mechanism to generate varied evaluation instances, enabling robust assessment of LLMs’ cross-calendar temporal reasoning.
We conduct comprehensive experiments on both open- and closed-source state-of-the-art LLMs over a range of dates spanning a $100$-year period from $1960$ to $2060$. We demonstrate that these LLMs achieve an average accuracy of only 34.534.5% across the evaluation dates, with none exceeding 8080% accuracy. We also investigate the impact of reasoning types, question formats, and temporal reasoning directions, identifying two key obstacles for LLMs: Future-Date Degradation, where LLMs struggle with cross-calendar temporal reasoning under future reference dates, and Calendar Asymmetry Bias, an asymmetry where reasoning from the Gregorian calendar is more accurate than the reverse. To advance LLMs’ cross-calendar temporal reasoning capabilities, we develop an LLM-powered Time Agent with tool-augmented code generation. Empirical results demonstrate that our method achieves an average accuracy of 95.3195.31%, surpassing competitive baselines. Our contributions are summarized as follows:
  • We introduce SPAN, a benchmark assessing LLMs’ cross-calendar temporal reasoning across ten directions, two reasoning types, and two question formats.
  • We propose a template-driven protocol for dynamic instance generation, enabling broad temporal evaluation while mitigating data contamination.
  • We conduct comprehensive experiments and in-depth analyses, revealing that LLMs struggle to reason across calendars and identifying the obstacles of Future-Date Degradation and Calendar Asymmetry Bias.
  • We develop an LLM-powered Time Agent using tool-augmented code generation, which achieves an average accuracy of 95.3195.31%, surpassing multiple strong baselines and demonstrating the potential of tool-augmented code generation for the cross-calendar temporal reasoning task.
Related Work
Temporal Reasoning Benchmarks for LLMs.
Temporal reasoning, the capability to perceive and understand the world’s dynamics, is essential for advancing LLMs toward artificial general intelligence (AGI). Early benchmarks such as TimeQA (chen2021datasetansweringtimesensitivequestions) and TimeDial (qin-etal-2021-timedial) concentrate on time-sensitive question answering and temporal commonsense in dialogue contexts. Recently, a new wave of benchmarks has emerged, targeting more fine-grained, complex temporal reasoning scenarios (chen2021datasetansweringtimesensitivequestions; liska2022streamingqa; kasai2023realtime; tan2023benchmarkingimprovingtemporalreasoning; wei-etal-2023-menatqa; tan-etal-2023-towards; wang2024trambenchmarkingtemporalreasoning; chu-etal-2024-timebench; wang-zhao-2024-tram; fatemi2024test; ge2025tremu; wei2025time; saxena2025lost). For example, TimeBench (chu-etal-2024-timebench) introduces a hierarchical benchmark for evaluating a wide spectrum of temporal reasoning capabilities, ranging from temporal relation classification to duration estimation and multi-hop time-aware question answering, spanning symbolic, commonsense, and event-centric tasks. StreamingQA (liska2022streamingqa) and RealtimeQA (kasai2023realtime) extend the focus to dynamic settings where new information arrives continuously, and LLMs are expected to revise their responses over time. Additionally, Lost in Time (saxena2025losttimeclockcalendar) explores how multimodal LLMs process time-related visual inputs, including analogue clocks and yearly calendars. These benchmarks collectively enable systematic evaluation of LLMs’ temporal reasoning across diverse tasks, complexities, and modalities.
Improving Temporal Reasoning Capabilities of LLMs.
In pursuit of human-level temporal reasoning, recent research has proposed various strategies to enhance LLMs (wei-etal-2023-menatqa; tan-etal-2023-towards; yang-etal-2023-upon; jain2023language; sutimo; xiong2024large; yang-etal-2024-enhancing-temporal). For instance, RemeMo (yang-etal-2023-upon) improves LLMs’ temporal understanding during pre-training by organizing events or sentences according to their temporal order, facilitating the model’s capture of complex temporal relationships. TG-LLM (xiong2024large) applies supervised fine-tuning on a synthetic temporal graph dataset, enabling the model to convert text into structured graphs for more effective temporal reasoning. Additionally, TempReason (yang-etal-2024-enhancing-temporal) leverages reinforcement learning to encourage the model to generate temporally coherent predictions, thereby improving the performance across diverse temporal reasoning tasks. These efforts advance the temporal reasoning capabilities of LLMs across various tasks.
SPAN: A Cross-Calendar Temporal Reasoning Benchmark for Large Language Models
In this section, we first formalize the cross-calendar temporal reasoning task. Then, we describe the data acquisition and template design, including cross-calendar data collection, a description of the newly developed cross-calendar conversion interface, and the construction of question–code template pairs. Finally, we present a novel evaluation protocol for dynamic instance generation.
Reasoning Type Question Format Question Template
Date- based Content question Today’s date on the $\{c_{s}\}$ calendar is $\{d^{r}_{c_{s}}\}$”. What was the $\{c_{t}\}$ calendar date $\{n_{d}\}$ days ago?
Today’s date on the $\{c_{s}\}$ calendar is $\{d^{r}_{c_{s}}\}$”. What is the $\{c_{t}\}$ calendar date $\{n_{d}\}$ days later?
Today’s date on the $\{c_{s}\}$ calendar is $\{d^{r}_{c_{s}}\}$”. What was the $\{c_{t}\}$ calendar date $\{n_{w}\}$ weeks ago?
Today’s date on the $\{c_{s}\}$ calendar is $\{d^{r}_{c_{s}}\}$”. What is the $\{c_{t}\}$ calendar date $\{n_{w}\}$ weeks later?
Polar question Today’s date on the $\{c_{s}\}$ calendar is $\{d^{r}_{c_{s}}\}$”. Was the $\{c_{t}\}$ calendar date $\{n_{d}\}$ days ago equivalent to the date $\{d^{e}_{c_{t}}\}$”?
Today’s date on the $\{c_{s}\}$ calendar is $\{d^{r}_{c_{s}}\}$”. Is the $\{c_{t}\}$ calendar date $\{n_{d}\}$ days later equivalent to the date $\{d^{e}_{c_{t}}\}$”?
Today’s date on the $\{c_{s}\}$ calendar is $\{d^{r}_{c_{s}}\}$”. Was the $\{c_{t}\}$ calendar date $\{n_{w}\}$ weeks ago equivalent to the date $\{d^{e}_{c_{t}}\}$”?
Today’s date on the $\{c_{s}\}$ calendar is $\{d^{r}_{c_{s}}\}$”. Is the $\{c_{t}\}$ calendar date $\{n_{w}\}$ weeks later equivalent to the date $\{d^{e}_{c_{t}}\}$”?
Festival- based Content question Today’s date on the $\{c_{s}\}$ calendar is $\{d^{r}_{c_{s}}\}$”. What was the $\{c_{t}\}$ calendar date of the $\{c_{s}\}$ festival $\{f_{c_{s}}\}$$\{n_{d}\}$ ago?
Today’s date on the $\{c_{s}\}$ calendar is $\{d^{r}_{c_{s}}\}$”. What is the $\{c_{t}\}$ calendar date of the $\{c_{s}\}$ festival $\{f_{c_{s}}\}$$\{n_{y}\}$ later?
Polar question Today’s date on the $\{c_{s}\}$ calendar is $\{d^{r}_{c_{s}}\}$”. Was the $\{c_{t}\}$ calendar date of the $\{c_{s}\}$ festival $\{f_{c_{s}}\}$$\{n_{y}\}$ years ago equivalent to the date $\{d^{e}_{c_{t}}\}$”?
Today’s date on the $\{c_{s}\}$ calendar is $\{d^{r}_{c_{s}}\}$”. Is the $\{c_{t}\}$ calendar date of the $\{c_{s}\}$ festival $\{f_{c_{s}}\}$$\{n_{y}\}$ years later equivalent to the date $\{d^{e}_{c_{t}}\}$”?
Table 1: The question templates. We use curly-brace placeholders to denote variables: $c_{s}$ and $c_{t}$ denote the source and target calendars; $f_{c_{s}}$ denotes the festival in the source calendar; $d^{r}_{c_{s}}$ denotes the reference date in the source calendar; $d^{e}_{c_{t}}$ denotes the expected date in the target calendar (used for polar questions); $n_{d}$, $n_{w}$, and $n_{y}$ denote temporal offsets in days, weeks, and years, respectively.
Figure 1: Overview of the proposed evaluation protocol. Given a user-specified Gregorian date as input, the process proceeds through four stages: ➀ Calendar Conversion. The Gregorian date is converted into its equivalents in five calendars via our search_calendar interface, yielding $(c_{s},d_{c_{s}}^{r},f_{c_{s}})$ pairs, with $d_{c_{s}}^{r}$ and $f_{c_{s}}$ denoting the reference date and the festival in the source calendar $c_{s}$, respectively. ➁ Template Matching. These pairs are further utilized to construct $(c_{s},d_{c_{s}}^{r},f_{c_{s}},c_{t})$ pairs. Here, $c_{s}$ and $c_{t}$ are selected specifically to ensure one is a Gregorian calendar and the other is a non-Gregorian calendar. These pairs are matched against all question–code template pairs to generate candidate pairs. ➂ Template Instantiation. For each candidate question-code template pair, we manually specify the remaining variables $(d_{c_{t}}^{e},n_{d},n_{w},n_{y})$. Afterwards, question–code pairs are generated by filling the template placeholders with all variables. ➃ Code Execution. Finally, we execute each code snippet to generate the gold answer.
Task Formulation
Typically, the cross-calendar temporal reasoning task involves both intra-calendar date reasoning and inter-calendar date conversion, requiring LLMs to understand temporal relationships within individual calendars and to grasp temporal alignments between different calendars. Formally, this process can be represented as follows:
${d_{c_{s}}^{r}}\xrightarrow{\texttt{Reasoning}}{d_{c_{s}}}\xrightarrow{\texttt{Converting}}{d_{c_{t}}^{t}}$ (1)
where $d_{c_{s}}^{r}$} and $d_{c_{s}}$} represent the reference date and reasoning date in the source calendar $c_{s}$}, respectively; $d_{c_{t}}^{t}$} denotes the target date derived by converting $d_{c_{s}}^{i}$} to the target calendar $c_{t}$}. Note that Equation (1) is symmetric, enabling the source and target calendars to be interchanged.
In this work, we focus on a practical and representative cross-calendar temporal reasoning scenario, namely, reasoning between the Gregorian calendar and other calendar systems, which is prevalent in many countries and regions (taqizadeh1939various; graumann2015problem). Accordingly, in our formulation, when either the source calendar $c_{s}$} or the target calendar $c_{t}$} is set to the Gregorian calendar, the other is set to a non-Gregorian calendar.
Data Collection and Template Design
Cross-Calendar Data Collection.
SPAN covers dates from three major calendar families: solar (i.e., the Gregorian, Persian and Shaka calendars), lunar (i.e., the Islamic calendar), and lunisolar (i.e., the Hebrew and Chinese lunar calendars). Calendar dates are collected using Python libraries and Wikipedia [^1] https://www.wikipedia.org. Specifically, the datetime library [^2] https://docs.python.org/3/library/datetime.html is employed to enumerate Gregorian dates, and these dates are subsequently converted to their equivalent Islamic, Hebrew, Shaka, and Persian calendar dates using the convertdate library [^3] https://pypi.org/project/convertdate, while the conversions to the Chinese lunar calendar are handled using the LunarCalendar library [^4] https://pypi.org/project/LunarCalendar. Additionally, festival dates for each calendar are sourced and curated from Wikipedia. Festivals are listed in Appendix A. Finally, the collected dates are aligned to construct cross-calendar entries, each containing equivalent dates across the aforementioned six calendars.
Cross-Calendar Conversion Interface.
We develop a unified interface, search_calendar, which converts a date or festival in the given calendar to its equivalents in multiple target calendars. The interface supports two parameter schemas: {calendar_name, year, month, day} to specify a date, and {calendar_name, year, festival_name} to specify a festival in the given calendar. Upon invocation, search_calendar returns a cross-calendar entry containing equivalent dates across six calendars, enabling subsequent temporal computations. In this work, we employ this interface in both the code templates and the proposed Time Agent. We provide a detailed description of search_calendar in Appendix D.
Question Template Design.
As shown in Table 1, we design a set of question templates, which are organized into two categories: (1) Date-Based Question Templates, which apply temporal offsets in days or weeks to a reference date and convert the resulting date to its equivalent in the target calendar. (2) Festival-Based Question Templates, which apply a year offset to a reference year, identify the corresponding festival date for the resulting year, and convert it to its equivalent in the target calendar. For each category, two question formats are included: polar questions (answered with “Yes” or “No”) and content questions (answered with specific dates). The question templates collectively define eight variables, as detailed below. $c_{s}$ and $c_{t}$ denote the source and target calendars; $d_{c_{s}}^{r}$ denotes the reference date in the source calendar; $d_{c_{t}}^{e}$ denotes the expected date in the target calendar (used for polar questions); $f_{c_{s}}$ denotes the festival in the source calendar; $n_{d}$, $n_{w}$, and $n_{y}$ denote temporal offsets in days, weeks, and years, respectively. Among these, $(c_{s}$, $c_{t}$, $d_{c_{s}}^{r}$, $f_{c_{s}})$ are determined during evaluation, whereas $(d_{c_{t}}^{e}$, $n_{d}$, $n_{w}$, $n_{y})$ are manually configurable.
Code Template Design.
For each question template, we design a corresponding code template in Python for answer generation. Code templates mirror the date-related variables defined in the corresponding question templates but differ in how these variables are represented. Question templates express variables in natural language, whereas code templates represent them as code snippets. For example, the natural language expression $\{n_{d}\}$ days later” corresponds to the code snippet “+ $\texttt{+ timedelta(days=}\{n_{d}\}\texttt{)}$”. We divide the code templates into two categories based on the reasoning direction: Gregorian-to-Others and Others-to-Gregorian, indicating conversions from the Gregorian calendar to other calendar systems and vice versa, respectively. For the Gregorian-to-Others category, we first use the datetime library to compute the reasoning date in the Gregorian calendar based on the given reference date, and then convert this result to the target calendar using our search_calendar interface. Conversely, for the Others-to-Gregorian category, we first convert the reference date to the Gregorian calendar using our search_calendar interface, and then perform further computations with the datetime library.
Evaluation Protocol for Dynamic Instance Generation
Building upon our prior data preparation, we propose a novel evaluation protocol for on-the-fly instance generation, as illustrated in Figure 1. This protocol takes a user-specified Gregorian date as input and produces a diverse set of instances for cross-calendar temporal reasoning. Specifically, this protocol consists of the following stages: ➀ Calendar Conversion. We utilize the search_calendar interface to convert the user-specific Gregorian date into its equivalents in the Chinese lunar, Islamic, Hebrew, Shaka, and Persian calendars. This yields a set of $(c_{s},d_{c_{s}}^{r},f_{c_{s}})$ pairs, where $d_{c_{s}}^{r}$ and $f_{c_{s}}$ denote the reference date and the festival in the source calendar $c_{s}$, respectively. ➁ Template Matching. Building on these pairs, we further construct $(c_{s},d_{c_{s}}^{r},f_{c_{s}},c_{t})$ pairs, where $c_{s}$ and $c_{t}$ are selected such that one is the Gregorian calendar and the other is a non-Gregorian calendar. These pairs are then matched with all question-code template pairs to yield candidate pairs for subsequent instantiation. ➂ Template Instantiation. For each question-code candidate template pair, we manually specify the remaining variables $(d_{c_{t}}^{e},n_{d},n_{w},n_{y})$ from a predefined set (see the Experimental Setup for details). These variables are then substituted into the corresponding placeholders in the question and code templates, producing paired questions and executable code. ➃ Code Execution. Finally, each code snippet is executed, and its output serves as the ground truth answer for the corresponding evaluation instance.
Experimental Setup
Evaluation Models.
We evaluate six open- and closed-source LLMs. Specifically, the open-source models include Llama-3.3-70B-Instruct (grattafiori2024llama), DeepSeek-V3-1226 (liu2024deepseek), and Qwen-2.5-72B-Instruct (yang2024qwen2), which are accessed via Hugging Face and deployed using the vLLM framework (kwon2023efficient); the closed-source models include GPT-4o (hurst2024gpt), Claude-3.7-Sonnet (anth2025claude), and Gemini-1.5-Pro (team2024gemini), which are accessed through their APIs. All models are evaluated using greedy decoding with a temperature of $0.0$ to ensure reproducibility.
Evaluation Metric.
To enable reliable and scalable evaluation, we adopt GPT-4o as an automatic evaluator. Given a question, a model-generated response, and the gold answer, GPT-4o is prompted to assess the correctness of the response. Accuracy is then computed over the entire test set based on GPT-4o’s judgments. The complete evaluation prompt is detailed in Appendix E, and we also present the agreement between the GPT-4o evaluator and human annotations in Appendix C.
Figure 2: Left: Accuracy of LLMs across evaluation dates ranging from July 11st, $1960$ to July 11st, $2060$ at five-year intervals (July 11st omitted for clarity). The average accuracy over time is annotated for each model. Right: Average output token counts of LLMs at sampled evaluation dates. To ensure comparability, model outputs are tokenized using OpenAI’s tiktoken tokenizer with the o200k_base encoding.
Evaluation Setup.
The variables ($c_{s}$, $d_{c_{s}}^{r}$, $f_{c_{s}}$ $c_{t}$) are dynamically determined by our evaluation protocol, while the remaining variables ($d_{c_{t}}^{e}$, $n_{d}$, $n_{w}$, $n_{y}$) are manually specified as follows. The offset variables $n_{d}$ and $n_{w}$ are set to integers in the range [$1$, $10$], and $n_{y}$ in the range [$1$, 55]. For polar questions, the variable $d_{c_{t}}^{e}$ is set to the answer date, ensuring that all polar questions yield “Yes” as the gold answer. Additionally, we set the user-specified Gregorian dates (evaluation dates) to July 11st of every fifth year from $1960$ to $2060$, resulting in $21$ evaluation dates over $100$ years. At each of the $21$ evaluation dates, 11,$780$ instances are generated, comprising $800$ date-based and $980$ festival-based questions, yielding a total of 3737,$380$ evaluation instances. Finally, all numeric month values in the generated instances are replaced with their corresponding textual month names according to their respective calendars.
Main Results
The left part of Figure 2 shows accuracy at five-year intervals from July 11st, $1960$ to July 11st, $2060$. The following conclusions are drawn based on empirical evidence:
Cross-Calendar Temporal Reasoning Remains Challenging. As shown in the right part of Figure 2, we observe that LLMs tend to produce lengthy outputs ($150\sim$$300$ tokens), indicating that they perform explicit reasoning rather than directly predicting dates for cross-calendar temporal questions. Despite these reasoning attempts, the average accuracy across LLMs and evaluation dates remains only 34.534.5%, with none exceeding 8080% accuracy across evaluation dates. This underscores substantial room for improvement in LLMs’ cross-calendar temporal reasoning capabilities.
Accuracy Stratification in Large Language Models. We find that the closed-source LLMs (e.g., GPT-4o and Claude-3.7-Sonnet) generally outperform the open-source ones (e.g., Llama-3.3-70B-Instruct and Qwen-2.5-72B-Instruct). Encouragingly, the open-source DeepSeek-V3 achieves accuracy on par with leading closed-source LLMs, whereas the closed-source Gemini-1.5-Pro demonstrates the poorest performance, with accuracy consistently below 3030% across evaluation dates. These observations reflect the strengths of commercial LLMs and progress in the open-source community.
Accuracy Exhibits Significant Temporal Dependency. Across all LLMs, accuracies tend to improve for past dates but drop sharply for future ones. This temporally-dependent pattern of accuracy also persists across different reasoning types, question formats, and temporal reasoning directions (see the Analysis section for details). These findings suggest that LLMs have a limited capability to perform cross-calendar temporal reasoning for future dates, a phenomenon we term Future-Date Degradation.
Experimental Results
Figure 3: Accuracy of date-based and festival-based cross-calendar temporal reasoning over the evaluation dates from July 11st, $1960$ to July 11st, $2060$ at five-year intervals (July 11st omitted for clarity). The average accuracy over time for each reasoning type is annotated.
Figure 4: Accuracy of content question and polar question over the evaluation dates from July 11st, $1960$ to July 11st, $2060$ at five-year intervals (July 11st omitted for clarity). The average accuracy over time for each question format is annotated.
Figure 5: Accuracy of Gregorian-to-Others and Others-to-Gregorian cross-calendar temporal reasoning over the evaluation dates from July 11st, $1960$ to July 11st, $2060$ at five-year intervals (July 11st omitted for clarity). The average accuracy over time for each group’s temporal reasoning directions is annotated.
Analysis
We empirically investigate the impact of various factors on LLMs’ cross-calendar temporal reasoning capabilities, including reasoning types, question formats, and temporal reasoning directions. All comparisons reported have undergone significance testing, with detailed results provided in Appendix C.
Date-Based vs. Festival-Based.
Figure 3 presents the accuracy of LLMs on two reasoning types. We observe that most LLMs perform better on festival-based reasoning, with gains between 2.872.87% and 12.6012.60%, except for Gemini-1.5-Pro and Qwen-2.5-72B-Instruct. This likely stems from the prevalence of festival dates in pretraining data, which reduces the difficulty of festival-based reasoning.
Polar Question vs. Content Question.
Figure 4 presents the accuracy of two question formats. We find that the accuracy on polar questions consistently exceeds that of content questions for all LLMs, with gains ranging from 3.133.13% to 37.0837.08% (18.8618.86% on average). This gap reflects the lower complexity of polar questions, whose answers are binary, resulting in a 5050% chance of correctness even without precise temporal reasoning.
Gregorian-to-Others vs. Others-to-Gregorian.
Figure 5 presents accuracy for two group of temporal reasoning directions: Gregorian-to-Others and Others-to-Gregorian. We see that all LLMs perform better on the Gregorian-to-Others group, with accuracy gains ranging from 3.973.97% to 17.4917.49%, particularly among higher-performing models like DeepSeek-V3 (17.4917.49%), GPT-4o (15.8215.82%), and Claude-3.7-Sonnet (15.7615.76%). We refer to this discrepancy as Calendar Asymmetry Bias in LLMs, conjecturing that it likely stems from the prevalence of Gregorian-origin expressions in pretraining data. Since modern web documents and textual resources predominantly use Gregorian timestamps as the primary temporal anchor, LLMs are more frequently exposed to conversions originating from the Gregorian calendar, while receiving limited exposure to the reverse direction.
Improving Cross-Calendar Temporal Reasoning of LLMs
In this section, we present a practical method to improve LLM performance on SPAN. Our experiments indicate that LLMs face challenges with direct reasoning in this task. Consequently, we rely on an external tool for date conversion and defer enhancing LLMs’ intrinsic reasoning to future work, emphasizing a pragmatic solution here.
Our Method.
We develop an LLM-powered Time Agent that combines LLM code-generation capabilities with our cross-calendar conversion interface search_calendar. In particular, it follows a three-step agentic workflow: ➀ GPT-4o is guided with the description of search_calendar in a few-shot prompting, enabling it to generate executable code snippets for solving the given question. ➁ The generated code snippet is executed via a code interpreter, and we obtain the execution results. ➂ The execution results are appended to the dialogue context as additional input, upon which GPT-4o produces the final answer. The complete prompts of Time Agent are provided in Appendix F.
Baselines.
We compare our method with the following competitive baselines:
  • Previous Best Results: The best results across all LLMs for each evaluation date, illustrated in Figure 2.
  • OpenAI-o1: A state-of-the-art closed-source reasoning model with high token consumption, requiring extensive reasoning chains to process questions.
  • GPT-4o w/ RAG: The GPT-4o model augmented with a retrieval-augmented generation (RAG) system, leveraging the Bing Search [^5] https://www.microsoft.com/en-us/bing for external retrieval.
  • GPT-4o: The GPT-4o model, which is used to assess the effects of introducing RAG.
Figure 6: Left: Accuracy of each method over the evaluation dates at five-year intervals from July 11st, $1960$ to July 11st, $2060$. Right: Average output token count of each method at sampled evaluation dates. Model outputs are tokenized using the tiktoken library with the o200k_base encoding.
Results.
As depicted in Figure 6, Time Agent demonstrates consistently high accuracy across all evaluation dates, achieving an average accuracy of 95.3195.31% while maintaining a minimal average output token count. The few errors in Time Agent primarily stem from occasional code execution failures or subtle logical bugs, which can be further mitigated through enhanced prompt engineering. By comparison, OpenAI-o1 attains the second-highest average accuracy of 59.2959.29%, though it remains significantly below that of Time Agent and incurs substantially greater token usage. Additionally, we observe that GPT-4o w/ RAG achieves an average accuracy of 43.6943.69%, and slightly outperforms GPT-4o on future evaluation dates, with an average improvement of merely 0.680.68%. However, this marginal gain requires generating more output tokens compared to GPT-4o, indicating that RAG provides negligible benefit for the cross-calendar temporal reasoning task. In summary, our results show that reasoning models and RAG are insufficient for the cross-calendar temporal reasoning task. By contrast, tool-augmented code generation provides a robust and efficient solution.
Conclusion
In this work, we introduce SPAN, a benchmark to evaluate the cross-calendar temporal reasoning capabilities of LLMs. Unlike prior benchmarks that are Gregorian-centric and time-invariant, SPAN enables cross-calendar temporal reasoning across ten reasoning directions, two reasoning types, and two question formats. To support evaluation at different temporal contexts, we design a novel template-driven protocol for time-variant evaluation. Based on this protocol, we perform extensive empirical evaluations on dates ranging from $1960$ to $2060$, revealing significant limitations in LLMs’ capability to reason across calendars, with the challenges including Future-Date Degradation and Calendar Asymmetry Bias. Furthermore, we demonstrate that an LLM-powered Time Agent with tool-augmented code generation achieves impressive performance on the cross-calendar temporal reasoning task.
In future work, we plan to extend SPAN to encompass a wider spectrum of calendar systems, enabling more comprehensive coverage of global temporal contexts. Moreover, we will adapt existing Gregorian-based temporal reasoning tasks (e.g., event ordering and duration estimation) to cross-calendar settings, facilitating a systematic evaluation of LLMs’ generalization capabilities and advancing their robustness in diverse temporal reasoning scenarios.
Acknowledgments
This work was independently conducted at Li Auto. We sincerely thank Li Auto for its generous support, which was essential to the successful completion of this work.
Appendix
A. Festival List
Table 2 lists a total of $21$ festivals from four calendar systems: Gregorian, Chinese lunar, Islamic, and Persian.
Calendar
Festival
Gregorian
calendar
- Halloween
- Christmas Day
- New Year’s Day
- Valentine’s Day
- International Women’s Day
- International Workers’ Day
- International Children’s Day
Chinese
lunar calendar
- Chinese New Year
- Lantern Festival
- Dragon Boat Festival
- Chinese Valentine’s Day
- Ghost Festival
- Mid-Autumn Festival
Islamic
calendar
- Hijri New Year
- Isra and Mi’raj
- Eid al-Fitr
- Eid al-Adha
Persian
calendar
- Persian New Year
- Sizdah Be-dar
- Mehregan Festival
- Tirgan Festival
Table 2: Festivals in four calendars.
B. Reliability of the GPT-4o Evaluator
To validate the reliability of the GPT-4o evaluator, we conduct a correlation test. Concretely, we sample $1,260$ examples from different LLMs and compare GPT-4o’s evaluations with human annotations using multiple correlation metrics.
As presented in Table 3, GPT-4o’s evaluations exhibit a high degree of concordance with human annotations, achieving strong metric scores ranging from $0.95$ to $0.98$, with consistently high agreement across individual LLM outputs. These results support GPT-4o’s reliability as an unbiased evaluator within our benchmark.
Model
Agreement
Accuracy
Spearman Kendall Cohen’s $\bm{\kappa}$
$0.98$
(12321232/$1260$)
$0.95$ $0.95$ $0.95$
GPT-4o
$0.98$
(205205/$210$)
$0.95$ $0.95$ $0.95$
Claude-3.7-
Sonnet
$0.97$
(203203/$210$)
$0.94$ $0.94$ $0.93$
Gemini-1.5-
Pro
$0.97$
(204204/$210$)
$0.88$ $0.88$ $0.87$
Llama-3.3-
70B-Instruct
$0.96$
(201201/$210$)
$0.88$ $0.88$ $0.88$
DeepSeek-V3
$1.00$
(210210/$210$)
$1.00$ $1.00$ $1.00$
Qwen-2.5-
72B-Instruct
$0.99$
(209209/$210$)
$0.99$ $0.99$ $0.99$
Table 3: Correlation metrics between the GPT-4o’s evaluations and human annotations.
C. Statistical Significance Testing
We conduct significance testing using $10,000$ bootstrap iterations to rigorously evaluate the robustness of our experimental findings. Tables 4, 5, and 6 present the results for the Festival-Based vs. Date-Based, Polar-Question vs. Content-Question, and Gregorian-to-Others vs. Others-to-Gregorian reasoning tasks, respectively. All pairwise comparisons reveal statistically significant differences with $p$ << $0.001$.
Model
Mean
Difference
95%Confidence
Interval
P-Value
GPT-4o
$0.16$ [$0.15$, $0.17$] $p$ << $0.001$
Claude-3.7-Sonnet
$0.16$ [$0.15$, $0.17$] $p$ << $0.001$
Gemini-1.5-Pro
$0.04$ [$0.03$, $0.05$] $p$ << $0.001$
Llama-3.3-70B-Instruct
$0.07$ [$0.06$, $0.07$] $p$ << $0.001$
Deepseek-V3
$0.17$ [$0.16$, $0.19$] $p$ << $0.001$
Qwen-2.5-72B-Instruct
$0.06$ [$0.05$, $0.07$] $p$ << $0.001$
Table 4: Significance test results ($10,000$ bootstrap iterations) for Festival-Based and Date-Based reasoning tasks.
Model
Mean
Difference
95%Confidence
Interval
P-Value
GPT-4o
$0.03$ [$0.02$, $0.04$] $p$ << $0.001$
Claude-3.7-Sonnet
$0.18$ [$0.17$, $0.18$] $p$ << $0.001$
Gemini-1.5-Pro
$0.11$ [$0.11$, $0.12$] $p$ << $0.001$
Llama-3.3-70B-Instruct
$0.22$ [$0.21$, $0.22$] $p$ << $0.001$
Deepseek-V3
$0.22$ [$0.21$, $0.23$] $p$ << $0.001$
Qwen-2.5-72B-Instruct
$0.37$ [$0.36$, $0.38$] $p$ << $0.001$
Table 5: Significance test results ($10,000$ bootstrap iterations) for Polar-Question and Content-Question reasoning tasks.
Model
Mean
Difference
95%Confidence
Interval
P-Value
GPT-4o
$0.16$ [$0.02$, $0.04$] $p$ << $0.001$
Claude-3.7-Sonnet
$0.16$ [$0.15$, $0.17$] $p$ << $0.001$
Gemini-1.5-Pro
$0.04$ [$0.03$, $0.05$] $p$ << $0.001$
Llama-3.3-70B-Instruct
$0.07$ [$0.06$, $0.07$] $p$ << $0.001$
Deepseek-V3
$0.17$ [$0.16$, $0.19$] $p$ << $0.001$
Qwen-2.5-72B-Instruct
$0.06$ [$0.05$, $0.07$] $p$ << $0.001$
Table 6: Significance test results ($10,000$ bootstrap iterations) for Gregorian-to-Others and Others-to-Gregorian reasoning tasks.
D. Description of search_calendar Interface
1Description
2--------
3An interface for cross-calendar date
4conversion. Given a calendar and either a
5specific date or a festival, it maps the
6given date to equivalent dates in multiple
7target calendar systems.
8
9Signature
10----------
11search_calendar(
12 calendar_name: str,
13 year: int,
14 month: Optional[int] = None,
15 day: Optional[int] = None,
16 festival_name: Optional[str] = None,
17 is_lunar_leap_month: bool = False
18) -> dict[str, str]
19
20Parameters
21----------
22- calendar_name : str
23The source calendar (e.g., Gregorian)
24
25- year: int
26Year in the source calendar. (e.g., 1998)
27
28- month: int, optional
29Month in the source calendar. (e.g., 12)
30Required when day is provided.
31
32- day: int, optional
33Day in the source calendar. (e.g., 23)
34Required when month is provided.
35
36- festival_name: str, optional
37Name of a festival. Required when neither
38month nor day is provided.
39
40- is_lunar_leap_month: bool, optional (
41default False)
42Whether the given lunar month is a leap
43month. Ignored for calendars that do not
44distinguish leap months.
45
46The interface supports two kinds of input:
47
48- {calendar_name, year, month, day[,
49is_lunar_leap_month]}: Specify a concrete
50date in the source calendar.
51
52- {calendar_name, year, festival_name}:
53Specify a festival in the source calendar.
54
55The parameter is_lunar_leap_month is
56relevant for lunisolar calendars (e.g.,
57Chinese Lunar calendar) where months can be
58leap months. It will be ignored for
59calendars that do not support this feature.
60
61Return
62-------
63- a cross-calendar entry: dict
64A dictionary mapping target calendars to
65their equivalent dates or festival dates for
66the given input.
67
68Examples
69--------
70>>> search_calendar("Gregorian", 2024, 6, 1)
71{’Gregorian’: 2024-06-01’, Chinese Lunar’:
722024-04-25’, ...}
73
74>>> search_calendar("Chinese Lunar", 2025,
75festival_name="Chinese New Year")
76{’Chinese Lunar’: 2025-01-29’, Gregorian’:
772025-01-29’, ...}
E. Evaluation Prompt
1# Role: Text Evaluation Specialist
2
3## Description
4You are a Text Evaluation Specialist, your
5role is to assess response accuracy
6objectively based on task guidelines.
7
8## Task
9
10### Objective
11Check if the response correctly answers
12the question using the answer‘.
13
14### Inputs
15Receive a tuple: (‘question‘, response‘,
16answer‘)
17- question‘: A date conversion query.
18- response‘: The given answer.
19- answer‘: The correct answer.
20
21### Evaluation Criteria
22- **Accuracy**: Verify if response matches
23answer‘.
24- **Ignore Formatting**: Disregard date
25format issues.
26
27## Task Output
28Use this JSON format for your evaluation, 1
29for correct, 0 for incorrect:
30‘‘‘json
31{
32 "accuracy": ""
33}
34‘‘‘
F. Prompts for Our Proposed Time Agent
1. Prompt for Code Generation
1# Role
2
3## Calendar System Date and Festival
4Converter
5You can accurately convert dates and
6festival observances across multiple
7calendar systems using the search_calendar
8function and functions from the datetime
9package.
10
11## Function Definition
12
13### Function: search_calendar
14
15### Description
16This function maps dates or festivals from a
17given calendar to their equivalents in
18multiple target calendars.
19
20### Input Parameters
21- calendar‘: which calendar to use
22[
23 "gregorian",
24 "lunar",
25 "islamic",
26 "hebrew",
27 "shaka",
28 "persian"
29]
30- year‘: the year number
31- month‘: the month number
32- day‘: the day number
33- festival‘: festival name
34- is_lunar_leap_month‘: set True if month
35is a lunar leap month
36
37### Usage Requirements
38Provide either:
391. The full date (year, month, and day), or
402. The year and a valid festival name.
41
42### Return
43A dictionary with the date in multiple
44calendars. Example:
45{
46 "islamic_date": "1369-3-13",
47 "lunar_date": "1949-11-15",
48 "hebrew_date": "5710-10-14",
49 "shaka_date": "1871-10-13",
50 "persian_date": "1328-10-13",
51 "gregorian_date": "1950-1-3"
52}
53
54## Supported Festivals
55
56### Gregorian Calendar
57[
58 "New Years Day",
59 "Valentines Day",
60 "International Womens Day",
61 "International Workers Day",
62 "International Childrens Day",
63 "Halloween", "Christmas Day"
64]
65
66### Chinese Lunar Calendar
67[
68 "Chinese New Year",
69 "Lantern Festival",
70 "Dragon Boat Festival",
71 "Chinese Valentines Day",
72 "Ghost Festival",
73 "Mid-Autumn Festival"
74]
75
76### Islamic Calendar
77[
78 "Hijri New Year",
79 "Isra and Miraj",
80 "Eid al-Fitr",
81 "Eid al-Adha"
82]
83
84### Persian Calendar
85[
86 "Persian New Year",
87 "Sizdah Be-dar",
88 "Mehregan Festival",
89 "Tirgan Festival"
90]
91
92## Month Name Mapping
93
94### Hebrew Calendar
95{
96 1: "Nisan",
97 2: "Iyyar",
98 3: "Sivan",
99 4: "Tammuz",
100 5: "Av",
101 6: "Elul",
102 7: "Tishri",
103 8: "Heshvan",
104 9: "Kislev",
105 10: "Teveth",
106 11: "Shevat",
107 12: "Adar (Adar I)",
108 13: "Veadar (Adar II)"
109}
110
111### Shaka Calendar
112{
113 1: "Chaitra",
114 2: "Vaishakha",
115 3: "Jyeshtha",
116 4: "Ashadha",
117 5: "Shravana",
118 6: "Bhadrapada",
119 7: "Ashwin",
120 8: "Kartika",
121 9: "Margashirsha",
122 10: "Pausha",
123 11: "Magha",
124 12: "Phalguna"
125}
126
127### Islamic Calendar
128{
129 1: "Muharram",
130 2: "Safar",
131 3: "Rabi al-Awwal",
132 4: "Rabi al-Thani (Rabi al-Akhir)",
133 5: "Jumada al-Awwal",
134 6: "Jumada al-Thani (Jumada al-Akhir)",
135 7: "Rajab",
136 8: "Shaban",
137 9: "Ramadan", 10: "Shawwal",
138 11: "Dhu al-Qidah",
139 12: "Dhu al-Hijjah"
140}
141
142### Persian Calendar
143{
144 1: "Farvardin",
145 2: "Ordibehesht",
146 3: "Khordad",
147 4: "Tir",
148 5: "Mordad",
149 6: "Shahrivar",
150 7: "Mehr",
151 8: "Aban",
152 9: "Azar",
153 10: "Dey",
154 11: "Bahman",
155 12: "Esfand"
156}
157
158# Workflow Guidance
159
160## You can convert from Gregorian calendar
161to other calendars
1621. Use the datetime package to process or
163calculate the Gregorian date.
1642. Use search_calendar with the Gregorian
165result (calendar="gregorian").
1663. Store the final answer in the variable
167answer on the last line. For date comparison
168, refer to Example 3 or 4.
169
170## You can convert from non-Gregorian
171calendars to Gregorian calendar
1721. Use search_calendar with the relevant
173non-Gregorian calendar and parameters.
1742. If you need further date calculations,
175extract "gregorian_date" and use the
176datetime package.
1773. Store the final answer in the variable
178answer on the last line. For date comparison
179, refer to Example 3 or 4.
180
181# Examples
182
183## Example1
184- Query: Todays date on the Gregorian
185calendar is "2020-7-1". What was the Islamic
186calendar date 7 days ago?
187
188- Code: ‘‘‘python
189from datetime import datetime, timedelta
190
191gregorian_date_today = datetime(2020, 7, 1)
192gregorian_date_7_days_ago = \
193gregorian_date_today - timedelta(days=7)
194
195calendar = search_calendar(
196 calendar="gregorian",
197 year=gregorian_date_7_days_ago.year,
198 month=gregorian_date_7_days_ago.month,
199 day=gregorian_date_7_days_ago.day
200)
201
202answer = calendar.get("islamic_date")
203‘‘‘
204
205## Example2
206- Query: Todays date on the Chinese lunar
207calendar is "2020-7-1". What is the
208Gregorian calendar date of the Chinese lunar
209event "Mid-Autumn festival" 5 years later?
210
211- Code: ‘‘‘python
212lunar_year = 2020 + 5
213
214calendar = search_calendar(
215 calendar="lunar",
216 year=lunar_year,
217 event="Mid-Autumn Festival"
218)
219
220answer = calendar.get("gregorian_date")
221‘‘‘
222
223## Example3
224- Query: Todays date on the Gregorian
225calendar is "1960-7-1". Was the Islamic
226calendar date 5 days ago equivalent to the
227date "1 Muharram 1380"?
228
229- Code: ‘‘‘python
230from datetime import datetime, timedelta
231
232gregorian_date_today = datetime(1960, 7, 1)
233gregorian_date_5_days_ago = \
234gregorian_date_today - timedelta(days=5)
235
236calendar = search_calendar(
237 calendar="gregorian",
238 year=gregorian_date_5_days_ago.year,
239 month=gregorian_date_5_days_ago.month,
240 day=gregorian_date_5_days_ago.day
241)
242
243year, month, day = map(int, calendar. \
244get("islamic_date").split("-"))
245
246answer = bool(
247 year==1380 and \
248 month == 1 and \
249 day == 1
250)
251‘‘‘
252
253## Example4
254- Query: Todays date on the Hebrew calendar
255is "1 Tammuz 5725". Was the Gregorian
256calendar date 3 days ago equivalent to the
257date "1965-6-28"?
258
259- Code: ‘‘‘python
260from datetime import datetime, timedelta
261calendar = search_calendar(
262 calendar="hebrew",
263 year=5725,
264 month=4,
265 day=1
266)
267
268gregorian_date_today = datetime.strptime( \
269calendar.get("gregorian_date"), "%Y-%m-%d")
270
271gregorian_date_3_days_ago = \
272gregorian_date_today - timedelta(days=3)
273
274answer = bool(
275gregorian_date_3_days_ago.year == 1965 and \
276gregorian_date_3_days_ago.month == 6 and \
277gregorian_date_3_days_ago.day == 28")
278‘‘‘
279
280# Output Restriction
281Only respond with a code block enclosed in
[…]
283any explanations or comments.
2. Prompt for Final Response Generation
1Given a question and its answer, where the
2answer is either a numeric date string or a
3boolean, generate a response as follows:
4
51. Identify the calendar mentioned in the
6question (e.g., Gregorian, Hebrew, Shaka,
7Islamic, or Persian).
8
92. If the calendar is Hebrew, Shaka, Islamic
10, or Persian, convert the numeric month in
11the answer to its corresponding month name
12using the mappings below and format the date
13as "{Month Name} {Day}, {Year}". For other
14calendars, keep the numeric date format.
15
163. If the answer is a boolean (True or False
17), return "Yes." or "No." respectively.
18
19Supported Calendars and Month Name Mappings:
20
21Hebrew Calendar:
22{
23 1: "Nisan",
24 2: "Iyyar",
25 3: "Sivan",
26 4: "Tammuz",
27 5: "Av",
28 6: "Elul",
29 7: "Tishri",
30 8: "Heshvan",
31 9: "Kislev",
32 10: "Teveth",
33 11: "Shevat",
34 12: "Adar (Adar I)",
35 13: "Veadar (Adar II)"
36}
37
38Shaka Calendar:
39{
40 1: "Chaitra",
41 2: "Vaishakha",
42 3: "Jyeshtha",
43 4: "Ashadha",
44 5: "Shravana",
45 6: "Bhadrapada",
46 7: "Ashwin",
47 8: "Kartika",
48 9: "Margashirsha",
49 10: "Pausha",
50 11: "Magha",
51 12: "Phalguna"
52}
53
54Islamic Calendar:
55{
56 1: "Muharram",
57 2: "Safar",
58 3: "Rabi al-Awwal",
59 4: "Rabi al-Thani (Rabi al-Akhir)",
60 5: "Jumada al-Awwal",
61 6: "Jumada al-Thani (Jumada al-Akhir)",
62 7: "Rajab", 8: "Shaban",
63 9: "Ramadan",
64 10: "Shawwal",
65 11: "Dhu al-Qidah",
66 12: "Dhu al-Hijjah"
67}
68
69Persian Calendar:
70{
71 1: "Farvardin",
72 2: "Ordibehesht",
73 3: "Khordad",
74 4: "Tir",
75 5: "Mordad",
76 6: "Shahrivar",
77 7: "Mehr",
78 8: "Aban",
79 9: "Azar",
80 10: "Dey",
81 11: "Bahman",
82 12: "Esfand"
83}
84
85Example1:
86- Query: Todays date on the Gregorian
87calendar is "2060-7-1". What was the Islamic
88calendar date 10 days ago?
89- Answer: 1483-1-22
90- Response: 22 Muharram, 1483
91
92Examples
93Example2:
94- Query: Todays date on the Islamic
95calendar is "2 Safar 1483". What was the
96Gregorian calendar date 1 days ago?
97- Answer: 2060-06-30
98- Response: 6 30, 2060
99
100Output Requirement:
101Provide only the final converted answer as a
102single English date string or "Yes"/"No"
103without additional explanations.

### v2

@@
 Introduction
-Recent advances in large language models (LLMs) have led to substantial progress across a range of reasoning tasks (qiao-etal-2023-reasoning; xu2024faithful; li2024understanding; fan-etal-2025-slam).
-Among these, temporal reasoning, as a fundamental capability for LLMs to understand real-world dynamics, has garnered increasing attention, leading to the development of multiple evaluation benchmarks (chen2021datasetansweringtimesensitivequestions; liska2022streamingqa; kasai2023realtime; tan2023benchmarkingimprovingtemporalreasoning; wei-etal-2023-menatqa; tan-etal-2023-towards; fatemi2024test; wang2024trambenchmarkingtemporalreasoning; chu-etal-2024-timebench; wang-zhao-2024-tram; ge2025tremu; wei2025time; saxena2025lost).
+Recent advances in large language models (LLMs) have led to substantial progress across a range of reasoning tasks (Qiao et al.
+2023; Xu et al.
+2024; Li et al.
+2024; Fan et al.
+2025).
+Among these, temporal reasoning, as a fundamental capability for LLMs to understand real-world dynamics, has garnered increasing attention, leading to the development of multiple evaluation benchmarks (Chen et al.
+2021; Liska et al.
+2022; Kasai et al.
+2023; Tan et al.
+2023b; Wei et al.
+2023; Tan et al.
+2023a; Fatemi et al.
+2024; Wang and Zhao 2024b; Chu et al.
+2024; Wang and Zhao 2024a; Ge et al.
+2025; Wei et al.
+2025; Saxena et al.
+2025a).
 Nevertheless, most existing efforts focus on temporal reasoning within the Gregorian calendar, neglecting the exploration of LLMs’ capability to reason across calendar systems.
-In practice, numerous countries and regions employ an alternative calendar system alongside the widely-used Gregorian calendar system (taqizadeh1939various; graumann2015problem).
+In practice, numerous countries and regions employ an alternative calendar system alongside the widely-used Gregorian calendar system (Taqizadeh 1939; Graumann 2015).
 For instance, both Saudi Arabia and China use the Gregorian calendar for official purposes, but Saudi Arabia follows the Islamic calendar for religious observances while China relies on the Chinese lunar calendar for traditional cultural activities.
@@
 The ground truth in certain temporal reasoning benchmarks is typically deterministic and publicly available through platforms such as Hugging Face and GitHub.
-Such accessibility may lead to evaluation data being absorbed into training corpora, potentially biasing evaluations (sainz-etal-2023-nlp; deng-etal-2024-investigating; balloccu2024leak).
+Such accessibility may lead to evaluation data being absorbed into training corpora, potentially biasing evaluations (Sainz et al.
+2023; Deng et al.
+2024; Balloccu et al.
+2024).
 In response to these challenges, we introduce SPAN, a croSs-calendar temPoral reAsoning beNchmark.
@@
 Temporal reasoning, the capability to perceive and understand the world’s dynamics, is essential for advancing LLMs toward artificial general intelligence (AGI).
-Early benchmarks such as TimeQA (chen2021datasetansweringtimesensitivequestions) and TimeDial (qin-etal-2021-timedial) concentrate on time-sensitive question answering and temporal commonsense in dialogue contexts.
-Recently, a new wave of benchmarks has emerged, targeting more fine-grained, complex temporal reasoning scenarios (chen2021datasetansweringtimesensitivequestions; liska2022streamingqa; kasai2023realtime; tan2023benchmarkingimprovingtemporalreasoning; wei-etal-2023-menatqa; tan-etal-2023-towards; wang2024trambenchmarkingtemporalreasoning; chu-etal-2024-timebench; wang-zhao-2024-tram; fatemi2024test; ge2025tremu; wei2025time; saxena2025lost).
-For example, TimeBench (chu-etal-2024-timebench) introduces a hierarchical benchmark for evaluating a wide spectrum of temporal reasoning capabilities, ranging from temporal relation classification to duration estimation and multi-hop time-aware question answering, spanning symbolic, commonsense, and event-centric tasks.
-StreamingQA (liska2022streamingqa) and RealtimeQA (kasai2023realtime) extend the focus to dynamic settings where new information arrives continuously, and LLMs are expected to revise their responses over time.
-Additionally, Lost in Time (saxena2025losttimeclockcalendar) explores how multimodal LLMs process time-related visual inputs, including analogue clocks and yearly calendars.
+Early benchmarks such as TimeQA (Chen et al.
+2021) and TimeDial (Qin et al.
+2021) concentrate on time-sensitive question answering and temporal commonsense in dialogue contexts.
+Recently, a new wave of benchmarks has emerged, targeting more fine-grained, complex temporal reasoning scenarios (Chen et al.
+2021; Liska et al.
+2022; Kasai et al.
+2023; Tan et al.
+2023b; Wei et al.
+2023; Tan et al.
+2023a; Wang and Zhao 2024b; Chu et al.
+2024; Wang and Zhao 2024a; Fatemi et al.
+2024; Ge et al.
+2025; Wei et al.
+2025; Saxena et al.
+2025a).
+For example, TimeBench (Chu et al.
+2024) introduces a hierarchical benchmark for evaluating a wide spectrum of temporal reasoning capabilities, ranging from temporal relation classification to duration estimation and multi-hop time-aware question answering, spanning symbolic, commonsense, and event-centric tasks.
+StreamingQA (Liska et al.
+2022) and RealtimeQA (Kasai et al.
+2023) extend the focus to dynamic settings where new information arrives continuously, and LLMs are expected to revise their responses over time.
+Additionally, Lost in Time (Saxena et al.
+2025b) explores how multimodal LLMs process time-related visual inputs, including analogue clocks and yearly calendars.
 These benchmarks collectively enable systematic evaluation of LLMs’ temporal reasoning across diverse tasks, complexities, and modalities.
 Improving Temporal Reasoning Capabilities of LLMs.
-In pursuit of human-level temporal reasoning, recent research has proposed various strategies to enhance LLMs (wei-etal-2023-menatqa; tan-etal-2023-towards; yang-etal-2023-upon; jain2023language; sutimo; xiong2024large; yang-etal-2024-enhancing-temporal).
-For instance, RemeMo (yang-etal-2023-upon) improves LLMs’ temporal understanding during pre-training by organizing events or sentences according to their temporal order, facilitating the model’s capture of complex temporal relationships.
-TG-LLM (xiong2024large) applies supervised fine-tuning on a synthetic temporal graph dataset, enabling the model to convert text into structured graphs for more effective temporal reasoning.
-Additionally, TempReason (yang-etal-2024-enhancing-temporal) leverages reinforcement learning to encourage the model to generate temporally coherent predictions, thereby improving the performance across diverse temporal reasoning tasks.
+In pursuit of human-level temporal reasoning, recent research has proposed various strategies to enhance LLMs (Wei et al.
+2023; Tan et al.
+2023a; Yang et al.
+2023; Jain et al.
+2023; Su et al.
+2024; Xiong et al.
+2024; Yang et al.
+2024b).
+For instance, RemeMo (Yang et al.
+2023) improves LLMs’ temporal understanding during pre-training by organizing events or sentences according to their temporal order, facilitating the model’s capture of complex temporal relationships.
+TG-LLM (Xiong et al.
+2024) applies supervised fine-tuning on a synthetic temporal graph dataset, enabling the model to convert text into structured graphs for more effective temporal reasoning.
+Additionally, TempReason (Yang et al.
+2024b) leverages reinforcement learning to encourage the model to generate temporally coherent predictions, thereby improving the performance across diverse temporal reasoning tasks.
 These efforts advance the temporal reasoning capabilities of LLMs across various tasks.
@@
 Note that Equation (1) is symmetric, enabling the source and target calendars to be interchanged.
-In this work, we focus on a practical and representative cross-calendar temporal reasoning scenario, namely, reasoning between the Gregorian calendar and other calendar systems, which is prevalent in many countries and regions (taqizadeh1939various; graumann2015problem).
+In this work, we focus on a practical and representative cross-calendar temporal reasoning scenario, namely, reasoning between the Gregorian calendar and other calendar systems, which is prevalent in many countries and regions (Taqizadeh 1939; Graumann 2015).
 Accordingly, in our formulation, when either the source calendar $c_{s}$} or the target calendar $c_{t}$} is set to the Gregorian calendar, the other is set to a non-Gregorian calendar.
@@
 We evaluate six open- and closed-source LLMs.
-Specifically, the open-source models include Llama-3.3-70B-Instruct (grattafiori2024llama), DeepSeek-V3-1226 (liu2024deepseek), and Qwen-2.5-72B-Instruct (yang2024qwen2), which are accessed via Hugging Face and deployed using the vLLM framework (kwon2023efficient); the closed-source models include GPT-4o (hurst2024gpt), Claude-3.7-Sonnet (anth2025claude), and Gemini-1.5-Pro (team2024gemini), which are accessed through their APIs.
+Specifically, the open-source models include Llama-3.3-70B-Instruct (Grattafiori et al.
+2024), DeepSeek-V3-1226 (Liu et al.
+2024), and Qwen-2.5-72B-Instruct (Yang et al.
+2024a), which are accessed via Hugging Face and deployed using the vLLM framework (Kwon et al.
+2023); the closed-source models include GPT-4o (Hurst et al.
+2024), Claude-3.7-Sonnet (Anthropic 2025), and Gemini-1.5-Pro (Team et al.
+2024), which are accessed through their APIs.
 All models are evaluated using greedy decoding with a temperature of $0.0$ to ensure reproducibility.
@@
 Table 2 lists a total of $21$ festivals from four calendar systems: Gregorian, Chinese lunar, Islamic, and Persian.
-Calendar
-Festival
-Gregorian
-calendar
-- Halloween
-- Christmas Day
-- New Year’s Day
-- Valentine’s Day
-- International Women’s Day
-- International Workers’ Day
-- International Children’s Day
-Chinese
-lunar calendar
-- Chinese New Year
-- Lantern Festival
-- Dragon Boat Festival
-- Chinese Valentine’s Day
-- Ghost Festival
-- Mid-Autumn Festival
-Islamic
-calendar
-- Hijri New Year
-- Isra and Mi’raj
-- Eid al-Fitr
-- Eid al-Adha
-Persian
-calendar
-- Persian New Year
-- Sizdah Be-dar
-- Mehregan Festival
-- Tirgan Festival
+Calendar Festival
+Gregorian calendar - Halloween - Christmas Day - New Year’s Day - Valentine’s Day - International Women’s Day - International Workers’ Day - International Children’s Day
+Chinese lunar calendar - Chinese New Year - Lantern Festival - Dragon Boat Festival - Chinese Valentine’s Day - Ghost Festival - Mid-Autumn Festival
+Islamic calendar - Hijri New Year - Isra and Mi’raj - Eid al-Fitr - Eid al-Adha
+Persian calendar - Persian New Year - Sizdah Be-dar - Mehregan Festival - Tirgan Festival
 Table 2: Festivals in four calendars.
@@
 These results support GPT-4o’s reliability as an unbiased evaluator within our benchmark.
-Model
-Agreement
-Accuracy
-Spearman Kendall Cohen’s $\bm{\kappa}$
-$0.98$
-(12321232/$1260$)
-$0.95$ $0.95$ $0.95$
-GPT-4o
-$0.98$
-(205205/$210$)
-$0.95$ $0.95$ $0.95$
-Claude-3.7-
-Sonnet
-$0.97$
-(203203/$210$)
-$0.94$ $0.94$ $0.93$
-Gemini-1.5-
-Pro
-$0.97$
-(204204/$210$)
-$0.88$ $0.88$ $0.87$
-Llama-3.3-
-70B-Instruct
-$0.96$
-(201201/$210$)
-$0.88$ $0.88$ $0.88$
-DeepSeek-V3
-$1.00$
-(210210/$210$)
-$1.00$ $1.00$ $1.00$
-Qwen-2.5-
-72B-Instruct
-$0.99$
-(209209/$210$)
-$0.99$ $0.99$ $0.99$
+Model Agreement Accuracy Spearman Kendall Cohen’s $\bm{\kappa}$
+- $0.98$ (12321232/$1260$) $0.95$ $0.95$ $0.95$
+GPT-4o $0.98$ (205205/$210$) $0.95$ $0.95$ $0.95$
+Claude-3.7- Sonnet $0.97$ (203203/$210$) $0.94$ $0.94$ $0.93$
+Gemini-1.5- Pro $0.97$ (204204/$210$) $0.88$ $0.88$ $0.87$
+Llama-3.3- 70B-Instruct $0.96$ (201201/$210$) $0.88$ $0.88$ $0.88$
+DeepSeek-V3 $1.00$ (210210/$210$) $1.00$ $1.00$ $1.00$
+Qwen-2.5- 72B-Instruct $0.99$ (209209/$210$) $0.99$ $0.99$ $0.99$
 Table 3: Correlation metrics between the GPT-4o’s evaluations and human annotations.
@@
 All pairwise comparisons reveal statistically significant differences with $p$ << $0.001$.
-Model
-Mean
-Difference
-95%Confidence
-Interval
-P-Value
-GPT-4o
-$0.16$ [$0.15$, $0.17$] $p$ << $0.001$
-Claude-3.7-Sonnet
-$0.16$ [$0.15$, $0.17$] $p$ << $0.001$
-Gemini-1.5-Pro
-$0.04$ [$0.03$, $0.05$] $p$ << $0.001$
-Llama-3.3-70B-Instruct
-$0.07$ [$0.06$, $0.07$] $p$ << $0.001$
-Deepseek-V3
-$0.17$ [$0.16$, $0.19$] $p$ << $0.001$
-Qwen-2.5-72B-Instruct
-$0.06$ [$0.05$, $0.07$] $p$ << $0.001$
+Model Mean Difference 95%Confidence Interval P-Value
+GPT-4o $0.16$ [$0.15$, $0.17$] $p$ << $0.001$
+Claude-3.7-Sonnet $0.16$ [$0.15$, $0.17$] $p$ << $0.001$
+Gemini-1.5-Pro $0.04$ [$0.03$, $0.05$] $p$ << $0.001$
+Llama-3.3-70B-Instruct $0.07$ [$0.06$, $0.07$] $p$ << $0.001$
+Deepseek-V3 $0.17$ [$0.16$, $0.19$] $p$ << $0.001$
+Qwen-2.5-72B-Instruct $0.06$ [$0.05$, $0.07$] $p$ << $0.001$
 Table 4: Significance test results ($10,000$ bootstrap iterations) for Festival-Based and Date-Based reasoning tasks.
-Model
-Mean
-Difference
-95%Confidence
-Interval
-P-Value
-GPT-4o
-$0.03$ [$0.02$, $0.04$] $p$ << $0.001$
-Claude-3.7-Sonnet
-$0.18$ [$0.17$, $0.18$] $p$ << $0.001$
-Gemini-1.5-Pro
-$0.11$ [$0.11$, $0.12$] $p$ << $0.001$
-Llama-3.3-70B-Instruct
-$0.22$ [$0.21$, $0.22$] $p$ << $0.001$
-Deepseek-V3
-$0.22$ [$0.21$, $0.23$] $p$ << $0.001$
-Qwen-2.5-72B-Instruct
-$0.37$ [$0.36$, $0.38$] $p$ << $0.001$
+Model Mean Difference 95%Confidence Interval P-Value
+GPT-4o $0.03$ [$0.02$, $0.04$] $p$ << $0.001$
+Claude-3.7-Sonnet $0.18$ [$0.17$, $0.18$] $p$ << $0.001$
+Gemini-1.5-Pro $0.11$ [$0.11$, $0.12$] $p$ << $0.001$
+Llama-3.3-70B-Instruct $0.22$ [$0.21$, $0.22$] $p$ << $0.001$
+Deepseek-V3 $0.22$ [$0.21$, $0.23$] $p$ << $0.001$
+Qwen-2.5-72B-Instruct $0.37$ [$0.36$, $0.38$] $p$ << $0.001$
 Table 5: Significance test results ($10,000$ bootstrap iterations) for Polar-Question and Content-Question reasoning tasks.
-Model
-Mean
-Difference
-95%Confidence
-Interval
-P-Value
-GPT-4o
-$0.16$ [$0.02$, $0.04$] $p$ << $0.001$
-Claude-3.7-Sonnet
-$0.16$ [$0.15$, $0.17$] $p$ << $0.001$
-Gemini-1.5-Pro
-$0.04$ [$0.03$, $0.05$] $p$ << $0.001$
-Llama-3.3-70B-Instruct
-$0.07$ [$0.06$, $0.07$] $p$ << $0.001$
-Deepseek-V3
-$0.17$ [$0.16$, $0.19$] $p$ << $0.001$
-Qwen-2.5-72B-Instruct
-$0.06$ [$0.05$, $0.07$] $p$ << $0.001$
+Model Mean Difference 95%Confidence Interval P-Value
+GPT-4o $0.16$ [$0.02$, $0.04$] $p$ << $0.001$
+Claude-3.7-Sonnet $0.16$ [$0.15$, $0.17$] $p$ << $0.001$
+Gemini-1.5-Pro $0.04$ [$0.03$, $0.05$] $p$ << $0.001$
+Llama-3.3-70B-Instruct $0.07$ [$0.06$, $0.07$] $p$ << $0.001$
+Deepseek-V3 $0.17$ [$0.16$, $0.19$] $p$ << $0.001$
+Qwen-2.5-72B-Instruct $0.06$ [$0.05$, $0.07$] $p$ << $0.001$
 Table 6: Significance test results ($10,000$ bootstrap iterations) for Gregorian-to-Others and Others-to-Gregorian reasoning tasks.